always, show all work and include units to get full credit. To clean the outside of your house you rent a small high-pressure water sprayer. The sprayers pump delivers slow-moving water at a pressure of 10,000,000 Pa (about 150 atmospheres). How fast can this water move if all of its pressure potential energy becomes kinetic energy as it flows through the nozzle of the sprayer? (Solution: 141, but what are the units?) 5-3. Engineers are trying to create artificial gravity in a ring-shaped space station by spinning it like a centrifuge. The ring is 100 m in radius. How quickly must the space station turn in order to give the astronauts inside it apparent weights equal to their real weights at the earths surface? (Solution: About 31, but what are the units?)
Answer #1

13) The question is misleading enough. 10,000,000 Pa is NOT 150 atm. It is around 98.7 atm

Now,

P = 1/2*rho*v2

98.7*1.01e5 * 2 = 1000*v2

v = 141 m/s

----------------------------------------------------------

5-3) mg = mv2 / r

cancel m on both sides, we have

g = v2 / r

v = sqrt (gr)

v = sqrt (9.8*100)

v = 31.1 m/s